# Hamilton-Jacobi as model reduction, extension to Newtonian particle mechanics, and a wave mechanical curiosity

### v1

@@ 1 Introduction
2 Hamilton-Jacobi systems for general Newtonian forces and an associated wave equation
Consider the Newtonian particle system with prescribed force function $f_{i}$, where the index $i=1,\ldots,3N$ ranges over the total number of position degrees of freedom with $N$ being the number of particles:
$\dot{v}_{i}=f_{i}(x,v,t)$ (1a)
@@ 2 Hamilton-Jacobi systems for general Newtonian forces and an associated wave equation
and allow forces $f$ for which $\partial_{x_{j}}f_{i}-\partial_{x_{i}}f_{j}$ does not necessarily vanish. Also, the force functions $f_{i}$ include the relevant mass of the particle of which $v_{i}$ is a velocity degree-of-freedom. Suppose now we consider a reduced set of trajectories satisfying the ansatz
$v_{i}(t)=G_{i}(x(t),t)$
for some functions $G_{i}$ to be determined.
Requiring the functions $G_{i}:\Omega\times(0,T)\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{3n}$ and $(0,T)\subset\mathbb{R}_{+}$ to satisfy the PDE
$\frac{\partial G_{i}}{\partial x_{j}}(x,t)\,G_{j}(x,t)+\frac{\partial G_{i}}{\partial t}(x,t)-f_{i}(x,G(x,t),t)=0,$ (2)
one notes that defining ‘reduced’ trajectories by integrating
$\dot{x}^{(r)}_{i}(t)=G_{i}\left(x^{(r)}(t),t\right);\qquad x^{(r)}(0)=x_{i}^{0},$ (3)
@@
$\dot{v}^{(r)}_{i}(t)=\ddot{x}^{(r)}_{i}(t)=f_{i}\left(x^{(r)}(t),G\left(x^{(r)}(t),t\right),t\right)$ (4)
and, in this sense, such reduced trajectories satisfy the system (1), but only for velocity initial conditions that satisfy
@@
and for $\left(x_{i}^{(r)}(t),t\right)\in\Omega\times(0,T)$. Of course, several functions, $G^{(I)}$, can be computed for different initial conditions on the velocities (indexed by $I$), making contact with the ideas presented in [sawant2006model] related to computing (overflowing) invariant manifolds of dynamical systems [^1] We note that due to the dependence on $t$ of the functions $G$, they do not correspond to representations of the standard notion of invariant manifolds, but the notion of ‘time-dependent’ invariant manifolds have also been used in the dynamical systems literature.. We will see that, by invoking a gradient ansatz for the function $G$, as described below, contact is established with the existence of more than one solution to the classical Hamilton-Jacobi equation for Hamiltonian systems, as discussed between (19)-(20) in Sec. 3.
Suppose now we further restrict the ansatz for the functions $G_{i}$ by demanding that they be gradients of a scalar valued function $S:\Omega\times(0,T)\to\mathbb{R}$:
@@
where we assume a common mass $m$ for all particles for notational simplicity. Then (2) becomes
@@
the functions $D_{i}$ characterize forces that do not arise as gradients, in the space of positions (configuration space), of a scalar potential and include the position-dependent, non-conservative ‘curl forces’ of [berry2015hamiltonian], as well as dissipative forces. We note that most velocity and time-dependent potentials $\tilde{F}$ result in non-conservative and/or dissipative forces. An example of a conservative velocity-dependent potential is as follows: if $v\in\mathbb{R}^{3N}$ is the array of velocity degrees-of-freedom of all $N$ particles and $\mathbb{I}$ is the identity matrix on $\mathbb{R}^{3N}$ and $A\in\mathbb{R}^{3N}$ (used below) is an array required on dimensional grounds, the force (i.e., the negative partial derivative w.r.t $x$) arising from the potential given by $x\cdot\left(\mathbb{I}-\frac{v}{|v|}\otimes\frac{v}{|v|}\right)A$ ($\Big$(in components $x_{i}\left(\delta_{ij}-\frac{1}{v_{k}v_{k}}v_{i}v_{j}\right)A_{j}\Big)$ is conservative, since it is orthogonal to the velocity at all instants and hence expends no power.
@@
and for $F$ of the uncoupled form $F(x,v,t)=F^{v}(v,t)+F^{x}(x,0)$ with $F^{v}(0,t)=0$, $\tilde{F}(x,v,t)=F^{v}(v,t)$ so that $\frac{\partial\tilde{F}}{\partial x_{i}}(x,v,t)=0$, (9) becomes
@@
We note that, formally, any suitably defined continuous solution to (9) or (10) defines, through (6) and (3), a reduced, in the sense of (3-4-5), set of generalized solutions (w.r.t regularity) for $x_{i}$ in (1) corresponding to the specific class of force functions defining (9-10).
@@
Moreover, from this perspective (9-10) may be considered as a consistent generalization of the classical Hamilton-Jacobi equation for conservative particle dynamics to include a dissipative and non-conservative force. However, with the ansatz (6) in force, it is natural to require $D_{i}$ in (10) to be a gradient for solutions to exist, and this leaves very few choices for combining $\left(x,\frac{1}{m}\nabla S(x,t),t\right)$ for the desired outcome (noting that $D_{i}$ cannot be a gradient of a scalar function in its first ($x$) argument by definition (8)); we discuss the case of general forces in Sec. 2.2). The simplest such choice is the one related to linear damping mentioned earlier, so that $D=-\frac{\nu}{m}\,\nabla S,\nu\geq 0$, resulting in
@@
up to an additive function of time which we disregard.
@@ 2.1 A wave equation corresponding to the scalar Hamilton-Jacobi equation (11)
Consider a complex-valued wave function $\Psi:\Omega\times[0,T]\to\mathbb{C}$,
@@
2.2 Hamilton-Jacobi system for general forces in Newtonian mechanics
@@
Substituting this form in (2), using (8) and defining $\tilde{D}_{i}:=D_{i}\left(x,\frac{1}{m}\nabla S+R,t\right)+\frac{\nu}{m}\partial_{i}S$, we obtain the following system of Hamilton-Jacobi equations for the fields $(S,R)$ corresponding to Newtonian mechanics of mass points under general force systems (using the notation $\partial_{i}=\partial_{x_{i}}$):
@@ 3 A derivation of the classical Hamilton-Jacobi equation for conservative particle mechanics
$\hat{S}[c,s]=\int_{u_{1}}^{u_{2}}\mathcal{L}\left(c(s(u)),s(u),c^{\prime}(s(u))\right)\,\frac{ds}{du}(u)\,du$ (18)
@@
Following the argument leading up to (20) we also have, using (19-20) and for $c^{\prime}(t)$ given by (21),
@@
(22) then shows that the $S^{(I)}$ function defined above (cf. discussion between (19-20)) satisfies the classical Hamilton-Jacobi equation
@@ Acknowledgment
I thank Janusz Ginster, Ambar Sengupta, and Dmitry Vorotnikov for reading this work and sharing their helpful comments.

### v2

@@ 1 Introduction
2 Hamilton-Jacobi equation, an associated wave equation, and H-J-like systems for general Newtonian forces
Consider the Newtonian particle system with prescribed force function $\hat{f}_{i}$, where the index $i=1,\ldots,3N$ ranges over the total number of position degrees of freedom with $N$ being the number of particles:
$\dot{v}_{i}={\color[rgb]{0,0,0}\frac{1}{m_{i}}\hat{f}_{i}(x,v,t)=:f_{i}(x,v,t)\qquad(\mbox{no sum on }i)}$ (1a)
@@
and allow (mass-normalized) forces $f$ for which $\partial_{x_{j}}f_{i}-\partial_{x_{i}}f_{j}$ does not necessarily vanish. Also, as shown, the force functions $f_{i}$ include the relevant mass of the particle of which $v_{i}$ is a velocity degree-of-freedom. Suppose now we consider a reduced set of trajectories satisfying the ansatz
$v_{i}(t)=G_{i}(x(t),t)$ (2)
for some functions $G_{i}$ to be determined. For a slow-fast system with $v$ as the fast variables and $x$ slow, and for which the fast dynamics approaches an equilibrium for fixed slow variables (the Tikhonov approach), such an ansatz would be natural without an explicit $t$ dependence in $G$.
To derive a set of governing equations for the functions $G_{i}$, we note that a necessary condition for the ansatz (2) to be valid is that it satisfy (1a). This implies that (3) below must hold. Conversely, let the functions $G_{i}:\Omega\times(0,T)\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{3n}$ and $(0,T)\subset\mathbb{R}_{+}$ satisfy the PDE
$\frac{\partial G_{i}}{\partial x_{j}}(x,t)\,G_{j}(x,t)+\frac{\partial G_{i}}{\partial t}(x,t)-f_{i}(x,G(x,t),t)=0.$ (3)
Then, defining ‘reduced’ trajectories by integrating
$\dot{x}^{(r)}_{i}(t)=G_{i}\left(x^{(r)}(t),t\right);\qquad x^{(r)}_{{\color[rgb]{0,0,0}i}}(0)=x_{i}^{0},$ (4)
@@
$\dot{v}^{(r)}_{i}(t)=\ddot{x}^{(r)}_{i}(t)=f_{i}\left(x^{(r)}(t),G\left(x^{(r)}(t),t\right),t\right).$ (5)
In this sense, such reduced trajectories satisfy the system (1), but only for velocity initial conditions that satisfy
@@
and for $\left(x_{i}^{(r)}(t),t\right)\in\Omega\times(0,T)$. Of course, several functions, $G^{(I)}$, can be computed for different initial conditions on the velocities (indexed by $I$), making contact with the ideas presented in [sawant2006model] related to computing (overflowing) invariant manifolds of dynamical systems [^2] We note that due to the dependence on $t$ of the functions $G$, they do not correspond to representations of the standard notion of invariant manifolds, but the notion of ‘time-dependent’ invariant manifolds has also been used in the dynamical systems literature.. The approach above for posing (3) is a straightforward and natural adaptation, to a time-dependent context, of methods for computing invariant manifolds for dynamical systems, see e.g., [sacker1965new, muncaster1983invariant], with [muncaster1983invariant] abstracting ideas from the Kinetic Theory of Gases [truesdell1980fundamentals] where the functions $G$ are called ‘Gross Determiners,’ with the important distinction from (2) that they do not explicitly depend on time. We will see that, by invoking a gradient ansatz for the function $G$, as described below, contact is established with the existence of more than one solution to the classical Hamilton-Jacobi equation for Hamiltonian systems, as discussed between (21)-(22) in Sec. 3.
Suppose now we further restrict the ansatz for the functions $G_{i}$ by demanding that they be components of the gradient of a scalar valued function $S:\Omega\times(0,T)\to\mathbb{R}$:
@@
where we assume a common mass $m$ for all particles for notational simplicity. The potential $S$ has physical dimensions of Action $(energy.time)$. Then (3) becomes
@@
We note here that the link between (3) with $f=0$ and the quadratic Hamilton-Jacobi equation is well-known and much studied in the literature of Optimal Transport, e.g., [villani2021topics]. Of course, our primary interest is in studying systems with general position and velocity dependent force systems.
@@
the chosen ‘decomposition’ of $f$ is motivated by the first term in (8) which is a gradient. Of course, a solely position-dependent $V$ is the most common type of force potential that arises in applications related to conservative systems.
The functions $D_{i}$ characterize forces that do not arise as gradients, in the space of positions (configuration space), of a scalar potential and include the position-dependent, non-conservative ‘curl forces’ of [berry2015hamiltonian], as well as dissipative forces. We note that most velocity and time-dependent potentials $\tilde{F}$ result in non-conservative and/or dissipative forces.
An example of a conservative velocity-dependent potential is as follows: if $v\in\mathbb{R}^{3N}$ is the array of velocity degrees-of-freedom of all $N$ particles and $\mathbb{I}$ is the identity matrix on $\mathbb{R}^{3N}$ and $A\in\mathbb{R}^{3N}$ (used below) is an array required on dimensional grounds, the force (i.e., the negative partial derivative w.r.t $x$) arising from the potential given by $x\cdot\left(\mathbb{I}-\frac{v}{|v|}\otimes\frac{v}{|v|}\right)A$ ($\Big$(in components $x_{i}\left(\delta_{ij}-\frac{1}{v_{k}v_{k}}v_{i}v_{j}\right)A_{j}\Big)$ is conservative, since it is orthogonal to the velocity at all instants and hence expends no power.
@@
and for $F$ of the uncoupled form $F(x,v,t)=F^{v}(v,t)+F^{x}(x,0)$ with $F^{v}(0,t)=0$, we have $V(x):=F(x,0,0)=F^{x}(x,0)$ and so $\tilde{F}(x,v,t)=F^{v}(v,t)$ so that $\frac{\partial\tilde{F}}{\partial x_{i}}(x,v,t)=0$, (10) becomes
@@
We note that, formally, any suitably defined continuous solution to (10) or (11) defines, through (7) and (4), a reduced, in the sense of (4-5-6), set of generalized solutions (w.r.t regularity) for $x_{i}$ in (1) corresponding to the specific class of force functions defining (10-11).
@@
Moreover, from this perspective (10-11) may be considered as a consistent generalization of the classical Hamilton-Jacobi equation for conservative particle dynamics to include a dissipative and non-conservative force. However, with the ansatz (7) in force, it is natural to require $D_{i}$ in (11) to be a gradient for solutions to exist, and this leaves very few choices for combining $\left(x,\frac{1}{m}\nabla S(x,t),t\right)$ for the desired outcome (noting that $D_{i}$ cannot be a gradient of a scalar function in its first ($x$) argument by definition (9)); we discuss the case of general forces in Sec. 2.2). The simplest such choice is the one related to linear damping mentioned earlier, so that $D=-\frac{\nu}{m}\,\nabla S,\nu\geq 0$, resulting in
@@
up to an additive function of time which we disregard. It is interesting to note that the development above provides a physical justification for including a viscosity regularization in the classical H-J equation whose solutions are known to develop shocks in $\nabla S$ from smooth initial conditions.
In order to consider particle systems with general, unequal masses, one consider (1a) written in the form
$M_{ij}\dot{{v}}_{j}=\hat{f}_{i}(x,v,t),$
where $M\in\mathbb{R}^{3N\times 3N}$ is a positive-definite diagonal matrix of masses. Then one can define
$\hat{v}_{i}:=M_{ik}v_{k}$
to write (1a) in the form
$\dot{\hat{v}}_{i}=\hat{f}_{i}\left(x,M^{-1}\hat{v},t\right).$
One now assumes the ansatze
$\hat{v}_{i}(x,t)=\hat{G}_{i}(x,t);\qquad\hat{G}_{i}:=\frac{\partial\hat{S}}{\partial x_{i}},$
which implies
[…]
We note that the physical dimensions of $\hat{S}$ is also that of $Action$ or $energy.time$. The corresponding analog of (3) is then given by
$\frac{\partial\hat{G}_{i}}{\partial x_{j}}(x,t)\,\hat{G}_{j}(x,t)+\frac{\partial\hat{G}_{i}}{\partial t}(x,t)-\hat{f}_{i}\left(x,M^{-1}\hat{G}(x,t),t\right)=0,$ (13)
and that of (8) as
$\frac{\partial}{\partial x_{i}}\left(\frac{1}{2}|\nabla\hat{S}|^{2}+\frac{\partial\hat{S}}{\partial t}\right)(x,t)-\hat{f}_{i}\left(x,M^{-1}\nabla\hat{S}(x,t),t\right)=0.$
The rest of the development follows in an obvious manner with appropriate adjustment for the single mass $m$ appearing in the earlier simplified development.
@@ 2.1 A wave equation corresponding to the scalar Hamilton-Jacobi equation (12)
We revert to the equal mass case. Consider a complex-valued wave function $\Psi:\Omega\times[0,T]\to\mathbb{C}$,
@@
In closing this section, we note that the Madelung Transformation (see, e.g.,[bush2020hydrodynamic, Sec. 2.2.1], [carles2012madelung]), converts the Schrödinger equation to a pair of equations, one of which is the Hamilton-Jacobi equation in the classical limit $h\to 0$ for a particle system without dissipation [leonard2012schrodinger, khesin2019geometry, von2012optimal]. Here, our main emphasis has been to derive a dissipative Hamilton-Jacobi equation from Newtonian mechanics of particle systems with linear damping. Furthermore, a ‘converse’ of the Madelung route is adopted to arrive at a dissipative Schrödinger equation from Newtonian mechanics under specific assumptions, via the latter’s dissipative Hamilton-Jacobi equation. More importantly, our primary result is (19) in the following Section for which there is no Schrödinger equation and a corresponding ‘Madelung route.’
2.2 Hamilton-Jacobi-like system for general forces in Newtonian mechanics
@@
Substituting this form in (3), using (9) and defining $\tilde{D}_{i}:=D_{i}\left(x,\frac{1}{m}\nabla S+R,t\right)+\frac{\nu}{m}\partial_{i}S$, we obtain the following system of Hamilton-Jacobi-like [^4] Strictly speaking, (19) is a nonlinear second-order quasilinear system. However, since the highest order derivative in $R$ is only first-order and noting the explicit and allowed nonlinearities involving $(\partial S,\partial R)$ in the second-line of (19), we feel it is justified to refer to (19) as a H-J-like system. equations for the fields $(S,R)$ corresponding to Newtonian mechanics of mass points under general force systems (using the notation $\partial_{i}=\partial_{x_{i}}$):
@@ 3 A derivation of the classical Hamilton-Jacobi equation for conservative particle mechanics
$\hat{S}[c,s]=\int_{u_{1}}^{u_{2}}\mathcal{L}\left(c(s(u)),s(u),c^{\prime}(s(u))\right)\,\frac{ds}{du}(u)\,du.$ (20)
@@
Following the argument leading up to (22) we also have, using (21-22) and for $c^{\prime}(t)$ given by (23),
@@
(24) then shows that the $S^{(I)}$ function defined above (cf. discussion between (21-22)) satisfies the classical Hamilton-Jacobi equation
@@ Acknowledgment
I thank Janusz Ginster, Ambar Sengupta, Dmitry Vorotnikov, and Johannes Zimmer for reading this work and sharing their helpful comments.